OpenStudy (anonymous):

an investor invested a total of 2000 in two mutual funds one fund earned an 8% profit while the other earned a 3% profit. if the investors total profit is $150 how much was invested in total on the two mutual funds. what is he equation i need to solve the problem?

OpenStudy (anonymous):

You invest X in the 7% fund and 2200-X in the 4% fund: .07X + .04(2200-X) = 115 .03X = 115-88 = =27 X = 27/.03 = 900 So, $900 invested @ 7% and $1300 @ 4%

OpenStudy (anonymous):

did that help

OpenStudy (anonymous):

yes, thank you!